The depth of water behind The Hoover Dam in Nevada is 165m. What is the water pressure at a depth of 165m? The weight density of water is 9800N/m2
1
Expert's answer
2012-09-27T10:58:52-0400
P = h * d * g
P = 165 * 9800 = 1617000 Pa = 1.617 MPa